OpenStudy (anonymous):

hat is the standard form equation of the line passing through (3, –2) with a slope of 3/5? a. 5x – 3y = –19 b. 5x – 3y = 19 c. 3x – 5y = 19 d. 3x – 5y = –19

OpenStudy (anonymous):

the percents of three types of tickets collected at the gate for a high school football game are shown. Ticket type: Student Adult Senior percent: 48% 28% 14% there is one other type of ticket that is not shown. it is a ticket for a child under 5.what percent of the tickets were of this type?

OpenStudy (anonymous):

the answer must be c or d; there is a theorem that indicates the slope 3/5 refers the 3 to the x and 5 to the y (one of the two is negative).

OpenStudy (anonymous):

overall you have two methods... method 1: use y=mx+b since you have the slope m and an x and a y from the point. substituting and solve for b: y=mx+b -2=(3/5)(3)+b b=-19/5 so y=(3/5)x-19/5 if you multiply through by 5 to get integers 5y=3x-19 rearrange this you get 19=3x-5y which is answer C

OpenStudy (anonymous):

method 2 would be to use point slope formula\[y-y _{1}=m(x-x _{1})\]where m is the slope and (x1, y1) is any point on the line. you just fill in the formula and put the line in standard form (the form that the answers are in).